Problem 5.8. For each of the following production functions, determine whether returns to scale are decreasing, constant, or increasing when capital and labor inputs are increased from K = L = 1 to K = L = 2. a. Q = 25K0.5L0.5 b. Q = 2K + 3L + 4KL c. Q = 100 + 3K + 2L d. Q = 5Ka Lb , where a+b=Β 


1
Expert's answer
2021-10-12T23:51:01-0400

Let's use "F(zK,zL)=AK^aL^b"

a. Q"=25K^{0.5}L^{0.5}"

Let Q"_0=F(K,L)=25K^{0.5}L^{0.5}" be initial production function then after multiplying by z we get:

"Q_1=F(zK,zL)=25(zK)^{0.5}(zL)^{0.5}=z^{0.5}z^{0.5}25K^{0.5}L^{0.5}=z^{1.0}Q_0"

In this case F "(zK,zL)>zF(K,L)"

This production function represents increasing returns to scale.

b. Q"=2K+3+4KL"

Let "Q_0=F(K,L)=2K+3L+4KL" be the initial production function. Let us multiply it by factor z and call it Q"_1"

"Q_1=F(zK,zL)=2(zK)+3(zL)+4(zK)(zL)\n=z(2K+3L+4zKL)"

So "F(zK,zL)>zF(K,L)"

Production function represents increasing returns to scale.

c. Let Q"_0=" F"(" K,L")=100K+3K+2L(" initial production function")"

After multiplying it by factor z:

Q"_1=F(zK,zL)=100+3K+2L"

In this case F"(zK,zL)<" "zF(K,L)"

"100+3(zK)+2(zL)<100z+3(zK)+2(zL)"

This is a decreasing returns to scale

d. "Q_0=F(K,L)=5K^aL^b" where "a+b=1"

Let's multiply it by factor z

"Q_1=F(zK,zL)=5(zK)^a(zL)^b=z^az^b5K^aL^b"

"z^{a+b}Q_0=zQ_0"

Here "F(zK,zL)=zF(K,L)"

The production function represents a constant returns state.